OpenStudy (anonymous):

In the following problem, which step uses the additive identity to get to the next step? 1. -18 + 73 + 18 2. -18 + 18 + 73 3. 0 + 73 4. 73 A. Step 1 B. Step 2 C. Step 3 D. Step 4

OpenStudy (automaticloveletter):

Step 3

OpenStudy (automaticloveletter):

The Additive Identity Property: The additive identity is 0 because a + 0 = a and 0 + a = a This is often written in one line... a + 0 = a = 0 + a Where a is any real number.

OpenStudy (anonymous):

thinks ; )

OpenStudy (automaticloveletter):

No problem ^_^